a) Calculate the equilibrium price and equilibrium quantity of renewable energy



resources.



b) Calculate price elasticity of supply using point elasticity method when renewable



energy sector is in equilibrium. Also, interpret the result.



c) What will happen to the equilibrium quantity and equilibrium price of renewable



energy resources if energy sector improves the technology? (Graph is not



required)



Qd = 20,000 - 3P



Qs = 15,000 + 2P



1
Expert's answer
2021-12-05T18:58:02-0500

a.

"Qd =Qs\\ 20,000 - 3P





= 15,000 + 2P\\20 000-15 000=5p\\5000=5p\\p=5000\\Q=15 000+2(10 000)\\=35 000"

b.

"=\frac{Change \space in\space quantity\space demanded}{Change\space in\space price} \times \frac{Initial Price}{Initial quantity}\\=2 \times \frac{10 000}{35 000}\\=0.57, elastic"

c.

equilibrium price will reduce because technology will lower the production cost .

equilibrium quantity will rise.